OpenStudy (anonymous):

Find four geometric means between 25 and 800.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 25 \times r^5=800}\) (you have to "jump"/"move" 5 times since knowing that you have FOUR missing means, 25 is the first term, and 800 is the 6th term.

OpenStudy (solomonzelman):

so, \(\normalsize\color{blue}{ 25 \times r^5=800}\) \(\normalsize\color{blue}{ r^5=800 \div 25}\) \(\normalsize\color{blue}{ r^5=32}\) \(\normalsize\color{blue}{ r^5=2^5}\) \(\normalsize\color{blue}{ r=2}\)

OpenStudy (anonymous):

ok thanks